Maya deposits
$4000
into an account that pays simple interest at a rate of
3%
per year. How much interest will she be paid in the first
6
years?

Answer by stanbon(75887)

Maya deposits
$4000
into an account that pays simple interest at a rate of
3%
per year. How much interest will she be paid in the first
6
years?
----
I = P*r*t
Interest = 4000*0.03*6 = 4000*0.18 = $720.00
------------
Cheers,
Stan H.
--------------